Julio is planning to save for his newborn daughter's college tuition, to be paid in 18 years. The amount he wishes to have by that time is $65,000. He plans to deposit $2,500 in a bank account every year for 18 years, starting next year. Suppose the account pays 2% interest on deposits, compounded annually. How much would he need to deposit right now, to make up the remaining amount necessary by the end of 18 years?
Answer #1

Future value of annuity=Annuity[(1+rate)^time period-1]/rate

=2500[(1.02)^18-1]/0.02

=2500*21.4123124

=$53530.781

Hence difference in amount needed at end of 18 years=65000-53530.781

=$11469.219

Hence amount to be deposited right now=11469.219*Present value of discounting factor(rate%,time period)

=11469.219/1.02^18

=11469.219*0.700159375

=$8030.28(Approx).